OpenStudy (anonymous):

solve 2 sec^2 x + tan^2 x - 3=0

OpenStudy (ash2326):

we have \[2 \sec^2 x+\tan ^2 x-3=0\] @Busterkitten what's the relation between tan^2 x and sec^2 x?

OpenStudy (ash2326):

Okay we have the relation as \[\tan^2 x+1=\sec^2x\] Remember it always it's an important relation:)

OpenStudy (ash2326):

We are given \[2sec^2x+tan^2x-3=0\] substitute \[tan^2 x=sec^2 x-1\] \[2 \sec^2 x+ \sec^2 x-1-3=0\] we get \[3 \sec^2x-4=0\] or \[\sec^2x= \frac{4}{3}\] we get \[\sec x= \pm \frac{2}{\sqrt 3}\] now when \[\sec x= \frac{2}{\sqrt 3}\] x= 30 degrees when \[\sec x= -\frac{2}{\sqrt 3}\] x= 360-30=330 so x is either 30 or 330 degrees

OpenStudy (anonymous):

wouldn't be also be 210 and 150 cause the radical made the the # - and +

OpenStudy (ash2326):

Yeah you are right. Missed them . Sorry:(

OpenStudy (anonymous):

kk I can understand and thank you